# Handling Label Noise via Instance-Level Difficulty Modeling and Dynamic Optimization

### v1

@@ Abstract
Recent studies indicate that deep neural networks degrade in generalization performance under noisy supervision. Existing methods focus on isolating clean subsets or correcting noisy labels, facing limitations such as high computational costs, heavy hyperparameter tuning process, and coarse-grained optimization. To address these challenges, we propose a novel two-stage noisy learning framework that enables instance-level optimization through a dynamically weighted loss function, avoiding hyperparameter tuning. To obtain stable and accurate information about noise modeling, we introduce a simple yet effective metric, termed wrong event, which dynamically models the cleanliness and difficulty of individual samples while maintaining computational costs. Our framework first collects wrong event information and builds a strong base model. Then we perform noise-robust training on the base model, using a probabilistic model to handle the wrong event information of samples. Experiments on five synthetic and real-world LNL benchmarks demonstrate our method surpasses state-of-the-art methods in performance, achieves a nearly 75% reduction in computational time and improves model scalability.
@@ 1 Introduction
Label noise is an unavoidable issue in training deep neural networks (DNNs), particularly when dealing with large-scale datasets annotated through web scraping [1], crowd-sourcing [2], or pre-trained models [3]. Apparently, label noise is likely to degrade the model performance [4], mainly because the model tends to fit the noise patterns. To be specific, recent studies have found that DNNs have the memorization effect [5], i.e., DNNs first learn clean patterns early in training, then start fitting noise later, and eventually overfit the noisy dataset.
@@
Figure 1: A bar chart is shown to illustrate the distribution of wrong events. From the perspective of noise, clean samples typically exhibit lower wrong event values, while noisy samples tend to have higher wrong event values. From the perspective of sample difficulty, samples located at the extremes of the distribution are far from the decision boundaries (e.g., cat and bird in the figure), exhibit typical features, and are classified with high confidence by the model, representing easy samples. In contrast, samples located in the central depression of the distribution are often near the decision boundaries of multiple classes (e.g., cat and leopard in the figure). These samples are classified with low confidence by the model and represent hard samples.
To alleviate this problem, researchers have explored the use of memorization effect for the learning with noisy labels (LNL) task, which can be classified into three types of approaches, i.e., sample selection, loss regularization and label correction. Sample selection [6, 7, 8] leverages early-stage loss to identify clean samples for training, but it totally neglects the value of noisy data [9]. To address this, loss regularization methods [10, 11] model both clean and noisy samples by incorporating regularization terms or designing robust loss functions. However, they struggle in more challenging scenarios [12] like instance-dependent noise because they don’t fundamentally correct the noisy labels. Recent state-of-the-art approaches, label correction methods partition datasets into different subsets, such as the clean/noise set [13, 14] or clean/hard/noise set [15, 12]. Then they preserve the given labels in clean set and generate pseudo labels for noise set to achieve remarkable improvement. However, current label correction methods have the following limitations.
Limitations. (1) Accurate partitioning strategies are always computationally intensive such as co-training frameworks [6, 14, 16], k-means clustering [17] and contrastive loss [16]. (2) Exhaustive hyperparameter tuning is necessary such as coefficients of loss terms [14, 10, 12], filtering threshold to split subsets  [6, 14, 16] and the warm-up period to obtain a model have not fitted the noise [9, 14, 12]. (3) Existing works [14, 15, 12] always assign identical coefficients to all samples in the same subset, neglecting individual differences in cleanliness and difficulty within each subset.
@@
Our Proposal. To address these limitations, we propose DODO, an instance-level Difficulty MOdeling and Dynamic Optimization framework to achieve robust training over noisy data. The key idea is to design a dynamically weighted loss function that captures both the cleanliness and difficulty of each individual samples, enabling instance-level optimization. The loss function consists of three terms, i.e., clean, noise, and hard components. The weights for clean and noise components are dynamically adjusted based on sample cleanliness, while the weight of hard component is determined by sample difficulty, 25indicating that model performance is highly correlated with the cleanliness and difficulty modeling.
To this end, we introduce a new noise-robust metric termed wrong event, which counts the frequency of mismatches between model predictions and given labels across all the epochs during training. This metric can effectively capture both the cleanliness and difficulty of each individual sample (see Fig.1), providing adaptive weights for the loss function. Notably, wrong event achieves high computational efficiency while enabling accurate and robust modeling at any training stage (see Fig.2). By aggregating information across the entire training process, wrong event demonstrates superior stability in modeling hard and noisy samples, effectively avoiding the inaccurate subset partitioning frequently happened in loss-based modeling methods due to the loss mutation of hard and noisy samples [18].
Overall, DODO leverages wrong event as a key metric in a two-stage process (see Fig.3). In the first stage, we adopt a conventional training strategy to collect wrong event information for each sample and obtain a competitive base model. In the second stage, we perform robust training on the basis of the base model, so the training effectiveness is improved. More specifically, a two-component beta mixture model (BMM) is employed to fit the distribution of wrong event, providing dynamic weights for the loss function. This iterative process ensures that wrong event information and the model performance mutually benefit each other, leading to much improvement throughout the training process.
We summarize our main contributions as follows:
  • We propose wrong event, a simple but effective metric that can reliably separate noisy samples from clean ones at any stage of training, even if the model is fitting noise. We provide detailed analysis to explain why wrong event works well.
  • We propose DODO, a two-stage noisy learning framework based on wrong event. Specifically, we first collect wrong event information and obtain a good base model. Then, we continue to train a more robust model with a dynamically weighted loss using wrong event information.
  • We run extensive experiments and show that DODO significant outperforms state-of-the-art methods by an average accuracy of 1.1% on synthetic and 0.8% on real-world datasets for LNL tasks, with a nearly 75% reduction in computational time, a better scalability to larger models compared to previous methods and without parameter tuning.
@@
Figure 2: A comparison between wrong event and loss. The baseline model is ResNet-18 trained on CIFAR-10 with 50% symmetric noise for 100 epochs. We show the loss distributions (the first row) and wrong event distribution (the second row) during training. The four columns, (a) (b) (c) (d), represent the distributions at epoch 20, 60, 100 and the entire training phase. In (d), the heavy lines represent the median value and the shaded areas are the interquartile ranges, respectively. Since wrong events are monotonically increasing based on historical statistics instead of current model prediction, when model overfits the dataset, wrong event values for all samples do not change, rather than converging to zero as loss values typically do. As a result, wrong event can clearly separate noisy data and clean data in all training stages, even if the model fits noise.
@@ 2 Related Works
Memorization & Forgetting. [4] observe that DNNs are sufficient for memorizing the entire dataset. [5] propose that DNNs follow the memorization effect. Researchers have studied the phenomenon and proposed some metrics. [19] proposed forgetting event, which counts how many times a model forgets a prediction of given labels. However, the model tends to predict ground truth labels rather than the given noisy labels for easy noisy samples consistently, resulting in the value of zero in early stage. [20] proposed first k-epoch learning, i.e., a sample has been predicted to its given label for consecutive $k$ epochs for the first time. However, $k$ is a hyperparameter which changes due to depending on the characteristics of the dataset. Though these metrics exhibit robustness during the model overfitting phase, their ability to distinguish between clean and noisy samples in the early stage is inferior to that of loss-based modeling. Our proposed wrong event metric demonstrates strong discriminative capabilities at all stages of model training, without long training periods.
@@
Learning with Noise Labels (LNL). LNL problem has been extensively explored in recent research [21, 12, 8, 22]. There are many variants of the LNL problem. Some methods [23, 24] suppose the existence of a small subset of clean data, while we do not rely on such assumptions. Facing this more challenging scenario, recent studies applied many techniques, such as co-training frameworks [6, 14, 16], k-means clustering [17], and contrastive loss [25, 16]. which require high computation and memory costs. To achieve instance-level optimization, dynamic loss functions are considered, but previous methods [9, 24] mainly focused on modeling clean and noisy samples dynamically, often overlooking the challenge of hard samples. Our simple yet robust wrong event modeling reduces computational costs, and the dynamic loss function simultaneously considers both sample cleanliness and difficulty.
@@
Pre-trained Models for LNL. Pre-trained models are recognized for their strong generalization performance with fast adaptation [26, 27]. Previous studies [21, 28] incorporate SimCLR [29] and MoCo [30] for label-noise learning, by using self-supervised pre-trained methods. Recent LNL methods [8, 22, 31] leverage pre-trained vision foundation models such as ResNet [32], ViT [26], CLIP [27] and ConvNeXt [33] to improve effectiveness and efficiency. However, larger models restricts the application of complex modeling approaches [8, 22], so they primarily rely on simple strategies such as the small-loss criterion or similarity threshold to filter out clean subsets for training, neglecting the value of noisy data. Our approach utilizes both clean and noisy information across the complete dataset, due to the high efficiency of wrong event. By doing so, DODO achieves superior model performance while keeping high efficiency.
3 Preliminary and Initial Findings
@@
In this section, we first introduce the problem setting of label noise learning. Subsequently, we formally define the concept of wrong event and present empirical evidence to demonstrate its robustness. Finally, we will discuss how to utilize the information provided by wrong event.
3.1 Settings of Noise Label Learning
@@
In the context of a $C$-class image classification problem, we denote the training dataset as $D_{\text{train}}=\{(x_{i},\bar{y}_{i})\}_{i=1}^{N}$, which consists of $N$ pairs of input images $x_{i}$ and their given labels $\bar{y}_{i}\in\{1,\dots,C\}$. In real-world scenarios, the given labels $\bar{y}_{i}$ may be corrupted due to various factors. We use $y_{i}$ to represent the ground truth label of $(x_{i},\bar{y}_{i})$, which remains inaccessible during the training phase. It is widely acknowledged that models trained on noisy labeled datasets $D_{\text{train}}$ using traditional cross-entropy loss often exhibit poor performance on clean test data. This limitation poses challenges when deploying models in real applications. Consequently, it becomes crucial to develop robust models capable of effectively handling datasets with noisy labels.
@@ 3.2 Observation of Wrong Event
We begin by formally defining the concept of wrong event. Consider a model $f(\cdot)$ with parameter $\theta$ trained over $T$ epochs. For a given sample $(x_{i},\bar{y}_{i})$, the wrong event is computed as follows:
@@
where $\mathbb{I}(\cdot)$ denotes the indicator function. This metric quantifies the cumulative number of epochs in which the model’s prediction $f(x_{i})$ disagrees with the given label $\bar{y}_{i}$. Based on empirical observations (See Fig.2), we found that wrong event can reliably separate noisy samples from clean samples at any stage of training, even if the model fits noise. We conduct extensive experiments to verify that wrong event is a more robust and informative metric than loss in more challenging noisy scenarios (see Appendix C) and provides more accurate modeling than other metrics (see Table 4).
Figure 3: Illustration of the proposed DODO framework. The training process is divided into two stages. In the first stage, prior knowledge, i.e., coarse distribution of wrong event, is obtained, and a base model that owns basic discrimination capability is captured. In the second stage, robust noise learning is performed. By using BMM, we obtain both cleanliness and difficulty information for individual samples, enabling instance-level dynamic optimization. The sample’s wrong event information and the model’s classification capability mutually benefit each other, leading to much improvement.
Before diving into the explanations behind this phenomenon, we need to make some definitions. Easy samples are those that lie far from the decision boundary, while hard samples are those located near the decision boundary. According to the widely accepted perspective in LNL [4, 34], the model should prioritize fitting samples in the following order: (1) easy clean samples, (2) hard samples (which can be either clean or noisy), and (3) easy noisy samples.
@@
We consider three explanations behind this phenomenon. 1) Prediction itself is a dynamic threshold. The model’s judgement can be likened to a dynamic threshold that adapts based on the characteristics of the data and the training progress. Unlike predefined thresholds that often require adjustments based on noise rates and noise types, wrong event leverages the model’s inherent learning dynamics, eliminating the need for hyperparameter tuning. To some extent, the wrong event maps the loss values to 0 or 1 based on a confidence-dependent dynamic threshold, where lower (higher) loss values are more likely to be mapped to 0 (1). 2) Wrong event is robust to noise. The variation of wrong event across consecutive epochs is significantly smaller compared to that of loss values. Specifically, while the loss of hard or noisy samples has been empirically observed to exhibit substantial fluctuations between two consecutive epochs [18], wrong event remains either unchanged or increases by at most 1 in the same interval. This stability stems from the fact that wrong event is a cumulative metric, reflecting the model’s historical prediction errors rather than instantaneous performance. Even if the model overfits to a noisy dataset, the outcome is simply that the wrong event values for all samples remain constant, rather than converging to zero as loss values typically do. 3) Wrong event reflects cleanliness and difficulty. Easy samples, which are located far from the decision boundary, exhibit consistent prediction behavior across consecutive epochs: clean(noisy) samples are consistently predicted correctly(incorrectly). As a result, these samples tend to occupy the extremes of the wrong event distribution. In contrast, hard samples, which lie close to the decision boundary, often experience fluctuating predictions, with their outcomes frequently flipping between similar classes. This instability causes hard samples to cluster in the middle region of the wrong event distribution, forming a low-lying trough. The distribution of wrong event provides valuable insights into both the cleanliness and the difficulty of each sample.
3.3 Utilize Wrong Event
In the context of LNL, probabilistic models are frequently employed to characterize the noise distribution. The two-component Gaussian mixture model (GMM) is commonly used to distinguish clean samples from noisy ones, and a threshold-based approach is employed to partition the dataset into different subsets [14, 16]. In all following equations, the symbol $w$ is used to denote the wrong event. The probability density function (pdf) of a mixture model of two components on wrong event is defined as:
@@ 3.3 Utilize Wrong Event
where $m_{k}$ are the mixing coefficients for the convex combination of each individual pdf $p(w|k)$. In our case, wrong event of easy samples, especially easy noisy samples, often concentrate at the distribution tails, leading to a monotonically increasing rather than unimodal distribution (see Fig.1,2). This phenomenon can be attributed to three main factors: 1) Hard samples [35, 36] often exhibit a long-tailed distribution in training; 2) Curriculum learning [37] suggests that models learn easy samples quickly but struggle with hard samples; 3) Training typically stops before the model fully memorizes the dataset, leading to an accumulation of easy noisy samples in the tail of the noise distribution.
These characteristics make Gaussian distribution models inadequate for fitting. To better model such distributions, we adopt a two-component beta mixture model (BMM), which is well-suited for capturing both symmetric, skewed and monotonically increasing distributions over $[0,1]$ due to the flexibility of the beta distribution [38]. The beta distribution over a (max) normalized wrong event $\in[0,1]$ is defined to have pdf:
$B(\alpha,\beta)=\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (3)
$p(w|\alpha,\beta)=\frac{w^{\alpha-1}\cdot(1-w)^{\beta-1}}{B(\alpha,\beta)}$ (4)
where $\alpha,\beta>0$ and $\Gamma(\cdot)$ is the Gamma function, and the mixture pdf is given by substituting the above into Eq. 2. We use an Expectation Maximization (EM) procedure to fit the BMM to the observations [9, 14]. Specifically, we introduce two variables:
@@
which defines the posterior probability of the given wrong event having been generated by mixture component $k$,
@@
which defines the cumulative distribution function (CDF) value of a beta mixture component $k$ for a given wrong event.
@@
where $\epsilon(w)$ represents the difficulty level of a sample with a given wrong event. The larger (smaller) $\epsilon(w)$ is, the closer (farther) the sample is to the decision boundary, indicating a higher (lower) difficulty level of the sample. It can be observed that when $w\to\text{min}\{w_{i}\}_{i=1}^{N}\ or\ \text{max}\{w_{i}\}_{i=1}^{N},\epsilon(w)\to 0$, and when $w\to\text{middle},\epsilon(w)\to 1$.
Due to the heterogeneous distribution of wrong event across different classes—where simpler classes typically exhibit lower wrong event means compared to harder ones—fitting a single distribution to the entire dataset proves insufficiently precise. To achieve more accurate modeling and maintain class balance, we establish $C$ BMM models to individually model the distribution of each class. For samples $i$ where $\bar{y_{i}}=c$, the posterior probabilities are derived from the corresponding BMM components $\mathcal{B}_{1}^{c}$ and $\mathcal{B}_{2}^{c}$, achieving more accurate and fine-grained modeling.

### v2

@@ Abstract
Recent studies indicate that deep neural networks degrade in generalization performance under label noise. Existing methods focus on data selection or label correction, facing limitations such as high computational costs, heavy hyperparameter tuning process, and coarse-grained optimization. To address these challenges, we propose a novel two-stage noisy learning framework that enables instance-level optimization through a dynamically weighted loss function, avoiding hyperparameter tuning. To obtain stable and accurate information about noise modeling, we introduce a simple yet effective metric, termed wrong event, which dynamically models the cleanliness and difficulty of individual samples while maintaining computational costs. Our framework first collects wrong event information and builds a strong base model. Then we perform noise-robust training on the base model, using a probabilistic model to handle the wrong event information of samples. Experiments on six synthetic and real-world LNL benchmarks demonstrate our method surpasses state-of-the-art methods in performance, achieves a nearly 75% reduction in storage and computational time, strongly improving model scalability. Our code is available at https://anonymous.4open.science/r/IDO-noisy_label/.
@@ 1 Introduction
Curating large training datasets through web scraping [1], crowd-sourcing [2], or pre-trained models [3] inevitably introduces label noise. It tends to degrade the performance of the trained deep neural network (DNN) model [4] due to DNN memorizing and thus overfitting the noisy samples [5].
To mitigate the negative impact of label noise, many works [6, 7, 8, 9, 10, 11, 12, 13, 14] consider the patterns of the clean and noisy samples shown in the DNN training process, i.e., DNNs first learn clean patterns early in training, then start fitting noise later, and eventually overfit the noisy dataset. However, one key problem of this strategy is that the noisy samples tend to show behavior similar to the clean but difficult samples – the clean examples that are close to the decision boundary of the model. To solve this problem, the most recent approaches [9, 11] try to distinguish noisy examples from hard examples by introducing the hardness of the samples into the loss function. However, these methods have the following limitations.
Limitations. (1) These methods are computationally intensive due to the overhead for measuring the hardness and cleanness of the samples and grouping them into different categories, e.g., clean, noisy, and difficult [15]; (2) these methods introduce extra hyperparameters, such as coefficients of loss terms, a cutoff threshold for grouping the samples, and a training epoch threshold to avoid the model fitting the noise, which are hard to tune [10, 12]; (3) these methods always assign identical coefficient to all samples within the same category, neglecting the differences of individual samples in both cleanliness and difficulty [16, 11].
@@
Figure 1: A bar chart illustrates the distribution of wrong events. In terms of sample cleanness, clean samples show lower wrong event, while noisy samples tend to have higher wrong event. In terms of sample difficulty, samples located at the extremes of the distribution represent easy samples, far from the decision boundaries (e.g., cat and bird in the figure). In contrast, samples located in the central depression of the distribution are often near the decision boundaries of similar classes, representing hard samples (e.g., cat and leopard in the figure).
@@
Our Proposal. To address these limitations, we propose IDO, an Instance-level Difficulty Modeling and Dynamic Optimization framework to achieve robust learning over noisy training data. Rather than relying on hyperparameters to regularize different terms in the loss function, IDO designs a dynamically weighted loss function that captures both the cleanliness and difficulty of each individual sample. This enables instance-level optimization without introducing any additional hyperparameters.
IDO makes this possible by proposing a new noise-robust metric to replace the classical loss-based metric which is known to be unstable and thus ineffective in distinguishing hard and noisy samples [17]. This metric, called wrong event, simply counts the frequency of mismatches between model predictions and the given labels, thus computationally efficient. Moreover, fitting a two-component beta mixture model to its distribution w.r.t. the training samples effectively measures the cleanliness and difficulty of each sample in a tuning-free manner (see Fig.1). Replacing the hard-coded coefficients of the loss terms with the corresponding measures that dynamically update epoch by epoch, IDO ensures accurate and robust modeling at any training stage (see Fig.2) yet achieves high computational efficiency.
We summarize our main contributions as follows: (1) We propose wrong event, a simple but effective metric that reliably separates noisy samples from clean ones at any stage of training, regardless of pretraining or not. We provide theoretical and experimental analysis to explain why wrong event works well whether the model has fitted noise. (2) We propose IDO that trains a robust model with a dynamically weighted loss using wrong event information where we are first to leverage probabilistic model to model data difficulty. (3) Our extensive experiments confirm that without hyperparameter tuning, IDO outperforms the state-of-the-art methods by an average accuracy of 1.6% on synthetic and 0.7% on real-world datasets for LNL tasks, with a 75% reduction in computational time, and better scalability to larger models compared to previous methods.
@@
Figure 2: A comparison between wrong event and loss. The baseline model is ResNet-18 trained on CIFAR-10 with 40% symmetric noise training for 100 epochs. We show the loss distributions (the first row) and wrong event distribution (the second row) during training. The four columns, (a) (b) (c) (d), represent the distributions at epoch 20, 60, 100 and the entire training phase. (a), (b), and (c) represent the early, middle, and late stages, respectively. In (d), the heavy lines represent the median value and the shaded areas are the interquartile ranges. Since wrong events are monotonically increasing based on historical statistics instead of current model prediction, when model overfits the dataset, wrong event values for all samples do not change, rather than converging to zero as loss values typically do. As a result, wrong event can clearly separate noisy data and clean data in all training stages, even if the model fits noise.
@@ 2 Related works
Memorization & Forgetting. [4] observe that DNNs are sufficient for memorizing the entire dataset. [5] shows that DNNs follow the memorization effect, i.e., clean samples and noisy samples exhibit distinct behaviors during the training process. Leveraging this observation, researchers have proposed other metrics to deal with noisy datasets. [18] proposed forgetting event, which counts how many times a model forgets a prediction of given labels. However, for easy noisy samples, the model tends to consistently predict ground truth labels rather than the given labels, resulting in the values of forgetting event all being zero in early stage. This makes it difficult to distinguish clean and noisy samples. [19] proposed first k-epoch learning (fkl), i.e., the time when a sample has been predicted to its given label for consecutive $k$ epochs for the first time. Samples with larger fkl are more likely to be noisy. However, $k$ is a hyperparameter that varies on different datasets. Although these metrics work well once the model training gets into the overfitting phase, they are less effective than loss in the early stage. Our proposed wrong event metric demonstrates strong discriminative capabilities regardless of the stages of model training (see Fig.2). More discussion about metrics is in Appendix C.3
@@
Learning with Noise Labels (LNL). LNL problem has been extensively explored in recent research [11, 13]. There are many variants of the LNL problem. Some methods [20, 21] assume the existence of a small subset of clean data, which our method does not require. In our more challenging scenario, which has to distinguish noisy and difficult samples, recent studies leverage techniques such as co-training frameworks [6, 16, 15], k-means clustering [22], and contrastive loss [23, 15], which unfortunately incur high computational and memory costs. Dynamic loss functions can be considered as instance-level optimization. However, previous methods [24, 21] mainly focus on modeling clean and noisy samples, while overlooking the challenge of modeling hard data, which our work targets.
@@
Pre-trained Models for LNL. Pre-trained models are known for their strong generalization ability [25, 26]. Previous studies [27, 28] learn the representations to avoid label noise, by using self-supervised pre-trained models such as SimCLR [29] and MoCo [30]. Recent LNL methods [12, 13, 14] leverage pre-trained vision models such as ResNet [31], ViT [25], CLIP [26] and ConvNeXt [32] to improve effectiveness and efficiency of LNL. However, large pre-trained models restrict the applicability of complex modeling approaches. Therefore, they [12, 13, 14] primarily rely on simple strategies such as the small-loss criterion or similarity threshold to discard the potential noisy data from training. This loses the opportunity to explore the valuable information hidden in these noisy data. Our approach is much more scalable because of the lightweight wrong event-based strategy. It is thus able to fully utilize both clean and noisy data, in turn achieving superior model performance.
3 Key ideas: wrong event and instance-level optimization
@@
In this section, we first introduce the problem of LNL. Subsequently, we define the concept of wrong event and present empirical evidence to demonstrate its robustness. Finally, we discuss how to fully leverage wrong event in LNL.
3.1 Noise label learning
@@
In the context of a $C$-class image classification problem , we denote a training dataset as $D_{\text{train}}=\{(x_{i},\bar{y}_{i})\}_{i=1}^{N}$, which consists of $N$ pairs of input images $x_{i}$ and their given labels $\bar{y}_{i}\in\{1,\dots,C\}$. In real-world scenarios, the given labels $\bar{y}_{i}$ may be corrupted due to various factors. We use $y_{i}$ to represent the ground truth label of $(x_{i},\bar{y}_{i})$, which remains inaccessible during training.
@@ 3.2 Observation of wrong event
We begin by formally defining the concept of wrong event. Consider a model $f(\cdot)$ trained over $T$ epochs. For a given sample $(x_{i},\bar{y}_{i})$, the wrong event is computed as follows:
@@
where $\mathbb{I}(\cdot)$ denotes the indicator function. This metric quantifies the cumulative number of epochs in which the prediction of the model $f(x_{i})$ disagrees with the given label $\bar{y}_{i}$. We observe that wrong event can reliably separate noisy samples from clean samples at any stage of training, even if the model has fitted noisy samples, as shown in Fig.2.
@@
In addition to the robustness of wrong event, i.e., the variance of wrong event across consecutive epochs is much smaller than that of loss values, where the high variance tends to make loss ineffective [17] in separating hard and noisy samples, wrong event effectively reflects both the cleanliness and difficulty of the training samples. Easy samples, which no matter clean or noisy, are located far from the decision boundary, exhibit consistent prediction behavior in consecutive epochs: clean (noisy) samples are consistently predicted correctly (incorrectly). As a result, these samples tend to occupy the extremes of the wrong event distribution. In contrast, hard samples, which lie close to the decision boundary, often experience fluctuating predictions, with their outcomes frequently flipping between similar classes. This instability causes hard samples to gather at the middle region of the wrong event distribution, forming a low-lying trough. Therefore, the distribution of wrong event offers valuable insights into both the cleanliness and the difficulty of each sample.
Our extensive theoretical and experimental analysis shows that wrong event is indeed a more robust and informative metric than loss with less variance, regardless of pretraining or not (see Appendix C) and provides more accurate measures than other metrics (see Table 4) in cleanliness and difficulty.
3.3 Measuring cleanliness and difficulty
To extract cleanliness and difficulty information from wrong event, we employ a two-component probabilistic model to fit the distribution of wrong event. In all the following equations, the symbol $w$ is used to denote the wrong event. The probability density function (pdf) of a two-component mixture model on wrong event is defined as:
@@ 3.3 Measuring cleanliness and difficulty
where $m_{k}$ are the mixing coefficients for the convex combination of each individual pdf $p(w|k)$.
Existing methods commonly use the Gaussian mixture model (GMM) to distinguish clean samples from noisy ones [16, 15]. However, in our case, the wrong event of easy samples, especially easy noisy samples, often focuses on the distribution tails, leading to a monotonically increasing rather than unimodal distribution (see Figs.1,2), making Gaussian distribution models inadequate for fitting. This phenomenon can be attributed to three main factors: 1) Hard samples [33, 34] often exhibit a long-tailed distribution in training; 2) Curriculum learning [35] suggests that models learn easy samples quickly but struggle with hard samples; 3) Training typically stops before the model fully memorizes the dataset, leading to an accumulation of easy noisy samples at the tail of the noise distribution.
To better model such distributions, we adopt a two-component beta mixture model (BMM), which is well suited to capture both symmetric, skewed, and monotonically increasing distributions due to the flexibility of the beta distribution [36]. We use an Expectation Maximization (EM) procedure [24, 16] to fit the BMM to the observations. Unlike previous method [24, 16], specifically, we introduce two more variables to fully use statistical information to build sample cleanliness and difficulty:
@@
which defines the posterior probability of a given $w$ having been generated by mixture component $k$,
@@
which defines the cumulative distribution function value of mixture component $k$ for a given $w$.
@@
where $\epsilon(w)$ represents the difficulty of a sample with the given wrong event. $\tau(w)$ reflects the posterior probability in $w$ belonging to clean or noisy distribution. $\lambda(w)$ measures how extreme $w$ is within its assigned distribution. Low $\epsilon(w)$ indicates $w$ lies in a high-confidence, typical region of one component, while high $\epsilon(w)$ suggests $w$ resides in ambiguous regions. The larger (smaller) $\epsilon(w)$ is, the closer (farther) the sample is to the decision boundary, indicating a higher (lower) difficulty level of the sample. We analyze the bounds of $\epsilon(w)$:
$10\leq\epsilon(w)\leq\tau_{1}(w)+\tau_{2}(w)=$ 1 (6)
Then we analyze its trend. When $w\approx\text{min}\{w_{i}\}_{i=1}^{N}$, we have $\tau_{1}(w)=1,r_{2}(w)=0,\lambda_{1}(w)=0,\lambda_{2}(w)=0$, then $\epsilon(w)\approx 1\times 0+0\times 1=0$, as does $w\approx\text{max}\{w_{i}\}_{i=1}^{N}$. When $w\to\text{middle}$, we have $\tau_{1}(w)\approx 0.5,r_{2}(w)\approx 0.5,\lambda_{1}(w)\approx 1,\lambda_{2}(w)\approx 0$, then $\epsilon(w)\approx 1$.
Due to the heterogeneous distribution of wrong event across different classes where simpler classes typically exhibit lower wrong event means compared to harder ones, fitting a single distribution to the entire dataset proves ineffective. To achieve more accurate modeling and maintain class balance, we establish $C$ BMM models to independently model the distribution of each class. For samples $i$ where $\bar{y_{i}}=c$, the posterior probabilities are derived from the corresponding BMM components $\mathcal{B}_{1}^{c}$ and $\mathcal{B}_{2}^{c}$, achieving more accurate and fine-grained modeling.